# Model Predictive Control of the Neural Manifold

### v1

@@ 4 Discussion
It may be important that there was a topological alignment between the latent spaces for neural activity and the stimulus set. The sVAE successfully discovered a dimensional mapping that separated the digits and their variants into 10 distinct and well-separated clusters (Figure 2A). There was also clear evidence of clustering in the neural latent space that mapped in an orderly way to which digit the stimulus was (Figure 3). Though it is beyond the scope of the present study, it would be interesting to explore how the stimulus latent spaces discovered by other dimensionality reduction methods impacts how the latent dynamics model performs in forecasting and control. It is also interesting to consider that the dynamics of the SNN and the latent mapping of the sVAE were both learned from the distribution of the stimulus, but through radically different methods. If topological alignment is required for control in this classification task and in other computational problems such as the ones explored by Susillo and Barak [29], it may speak to a simple but profound theory that derives from the ideas of James [45] and Hebb [46]: that learning is a process of aligning the latent dynamics of neural circuits to the latent dynamics of the physical world. Optimal control, both as theory and as a method for more precise experimental manipulation, may be of benefit in testing this theory in biological systems.

### v2

@@ 4 Discussion
It may be important that there was a topological alignment between the latent spaces for neural activity and the stimulus set. The sVAE successfully discovered a dimensional mapping that separated the digits and their variants into 10 distinct and well-separated clusters (Figure 2A). There was also clear evidence of clustering in the neural latent space that mapped in an orderly way to which digit the stimulus was (Figure 3). Though it is beyond the scope of the present study, it would be interesting to explore how the stimulus latent spaces discovered by other dimensionality reduction methods impacts how the LDM performs in forecasting and control. Similarly, alignment has been explored in relating the neural manifolds between organisms on the same tasks. Future work should explore if a controller trained one on organism’s manifold can be transferred to another by aligning these latent representations [59, 60]. If topological alignment is required for control in this classification task and in other computational problems such as the ones explored by Susillo and Barak [30], it may speak to a simple but profound theory that derives from the ideas of James [61] and Hebb [62]: that learning is a process of aligning the latent dynamics of neural circuits to the latent dynamics of the physical world. MPC, both as theory and as a method for more precise experimental manipulation, may be of benefit in testing this theory in biological systems.
5 Appendix
| Trial Length (ms) | Learning Rate | Training Epochs | Batch Size | ||
| 0.5 | 0.99 | 50 | 33 | 128 |
| Layer | Output Size | Kernel Size | Stride | Padding | Activation |
| Encoder | |||||
| Input | 28 x 28 | – | – | – | – |
| Conv Layer 1 | 32 x 28 x 28 | 3 x 3 | 1 | 1 | ReLU |
| Conv Layer 2 | 64 x 14 x 14 | 3 x 3 | 2 | 1 | ReLU |
| Conv Layer 3 | 64 x 7 x 7 | 3 x 3 | 2 | 1 | ReLU |
| Conv Layer 4 | 64 x 7 x 7 | 3 x 3 | 1 | 1 | ReLU |
| Flatten | 3136 | – | – | – | – |
| Linear Layer 1 | 256 | – | – | – | ReLU |
| Latent Mean () | 2 | – | – | – | – |
| Latent Log-Variance () | 2 | – | – | – | – |
| Decoder | |||||
| Linear Layer 1 | 256 | – | – | – | ReLU |
| Linear Layer 2 | 3136 | – | – | – | ReLU |
| Unflatten | 64 x 7 x 7 | – | – | – | – |
| ConvTranspose Layer 1 | 64 x 7 x 7 | 3 x 3 | 1 | 1 | ReLU |
| ConvTranspose Layer 2 | 64 x 13 x 13 | 3 x 3 | 2 | 1 | ReLU |
| ConvTranspose Layer 3 | 32 x 27 x 27 | 3 x 3 | 2 | 0 | ReLU |
| ConvTranspose Layer 4 | 29 x 29 | 3 x 3 | 1 | 0 | Sigmoid |
The output of the final deconvolutional layer was trimmed to produce a dimensional output image. This architecture was adapted from the example found in https://github.com/rasbt/stat453-deep-learning-ss21/blob/main/L17/1_VAE_mnist_sigmoid_mse.ipynb.
| Learning Rate | Batch Size | Training Epochs | |
| 128 | 57 | 0.5 |
| Layer | Dimensionality | BatchNorm | Activation |
| Encoder | |||
| Input | 500 | – | – |
| Hidden Layer 1 | 100 | ✓ | ReLU |
| Hidden Layer 2 | 50 | ✓ | ReLU |
| Hidden Layer 3 | 10 | ✓ | ReLU |
| Latent Mean () | 2 | – | – |
| Latent Log-Variance () | 2 | – | – |
| Decoder | |||
| Hidden Layer 1 | 10 | ✓ | ReLU |
| Hidden Layer 2 | 50 | ✓ | ReLU |
| Hidden Layer 3 | 100 | ✓ | ReLU |
| Hidden Layer 4 | 500 | ✓ | ReLU |
| Output | 500 | – | Sigmoid |
| Learning Rate | Batch Size | Training Epochs | |
| 256 | 20 | 0.001 |
| Learning Rate | Batch Size | Training Epochs | |
| 256 | 128 | 0.001 |
Data Availability
All code used for simulation, model fitting, control, and figure generation can be found at: https://github.com/melizalab/Neural-Manifold-MPC.git.
Acknowledgements
This work was supported by National Institutes of Health grant 1R01-DC018621, National Science Foundation grant IOS-1942480, and the University of Virginia Brain Institute.